# Singularity-free cosmology from interactions in the dark sector

### v1

@@ V.2.3 Four Einstein Fixed Points
Fig.s 15 - 17 show a first integral surface in the 3-D phase space, corresponding to a trajectory in Fig. 9 which intersects the $a^{\prime\prime}=0$ curve four times. Fig. 15 shows trajectories with positive spatial curvature, which all have a bounce. One CFS is plotted, which passes through one of the Einstein fixed points, which is a saddle, and loops round the Einstein fixed point above it, which is a center. A subset of trajectories within this CFS are cyclic, and have an initial accelerated expansion followed by a decelerated period, but no late-time acceleration. Bouncing trajectories within this CFS have one decelerated period, and those outside the CFS have two decelerated periods, both with a late-time acceleration. These models all contract from $dS_{2-}$, then accelerate through a bounce and become flatter as they start to expand, with those outside the CFS bouncing during a quasi de-Sitter phase; they then have one or two decelerated expanding phases, followed by a late-time acceleration as they evolve to $dS_{2+}$, therefore these bouncing models can all qualitatively match the observed Universe. The Einstein fixed point close to $X=1$ is a center, and the Einstein fixed point close to $X=0$ is a saddle. We expect this saddle fixed point to form part of another CFS, which would loop around the center Einstein fixed point close to $X=1$, however we do not plot this separatrix as the numerics breaks down at high energy. We do not expect trajectories within this CFS to be qualitatively interesting; we expect bouncing trajectories to always accelerate, and cyclic trajectories to have no late-time acceleration. Fig. 16 shows the view of the 3-D phase space through the $Z$-$X$ plane, where it is clearer which trajectory this surface corresponds to in the 2-D phase space in Fig. 9.
@@
Figure 15: The $X$-$Y$-$Z$ phase space corresponding to Fig. 9, where $w_{x}=-0.95$ and $q=3.1$. The trajectories themselves are plotted on a first integral surface where $X_{0}=0.14$ and $Z_{0}=0.3$, which corresponds to a trajectory in Fig. 9 that intersects the red zero acceleration curve four times. In this figure, only the trajectories with positive spatial curvature are plotted. Four Einstein fixed points (cyan) exist at $Y=0$ on the zero acceleration surface. The only trajectories that are not viable are the cyclic trajectories within the CFS, as during expansion there is no late-time acceleration. All other trajectories bounce once and contract from $dS_{2-}$, then expand towards $dS_{2+}$. Those within the CFS have one decelerated period followed by late-time acceleration, and those outside have two decelerated phases with a late-time acceleration.
@@
Figure 16: The side view of Fig. 15. The cyclic trajectories cross the red zero acceleration surface once during expansion. All other trajectories intersect this surface two or four times during expansion, accelerating when they are above the surface and decelerating when they are below it.
@@
Fig. 17 shows the expanding flat and open trajectories. These emerge from $dS_{1+}$ and $dS_{4+}$, respectively, and intersect the $a^{\prime\prime}=0$ surface four times. This means flat and open models have two periods of deceleration and three accelerated phases, including a late-time acceleration as they expand towards $dS_{2+}$.
@@
Figure 17: The expanding flat (green) and open (blue) trajectories in the $X$-$Y$-$Z$ phase space corresponding to Fig. 9, where $w_{x}=-0.95$ and $q=3.1$. The trajectories themselves are plotted on a first integral surface where $X_{0}=0.14$ and $Z_{0}=0.3$, which corresponds to a trajectory in Fig. 9 which intersects the red zero acceleration curve four times. The flat and open trajectories emerge from $dS_{1+}$ and $dS_{4+}$, respectively, and cross the red zero acceleration surface four times, eventually expanding toward $dS_{2+}$ at low energy.
@@ VI Conclusions
In this paper we have studied the dynamics of FLRW models with standard pressureless dark matter interacting with dark energy with a quadratic equation of state (EoS) (2). This EoS introduces two energy scales: $\rho_{\Lambda}$ turns out to be an effective cosmological constant, which plays the role of the asymptotic state of all models evolving toward low energies, and $\rho_{*}>\rho_{\Lambda}$ characterises the quadratic term. The interaction is also quadratic and characterised by an energy scale $\rho_{i}$, and defined so that energy flows from the dark energy to the dark matter. Given this form of interaction, the dark matter energy density $\rho_{m}$ is always positive, while the dark energy density $\rho_{x}$ remains positive if it was always positive in the past.
@@
This work is an extension of Paper I Burkmar and Bruni (2023), where we studied the dynamics of FLRW models containing dark matter, radiation and dark energy with a quadratic EoS, without any interaction term. In Paper I Burkmar and Bruni (2023), we found non-singular models that had an expanding decelerated phase and a late-time acceleration, however the decelerated period was lost when the effective cosmological constants were set to realistic energy scales, i.e. when $10^{-120}<\mathcal{R}=\rho_{\Lambda}/\rho_{*}<10^{-60}$. Here, our aim was two-fold. First, to see if including an interaction between the dark components allows for non-singular models, and second, to see if these non-singular models could evolve over realistic energy scales, with a decelerated period and late-time accelerated expansion.
@@
In Sec. II we presented the system of equations both in 2-D, where we take $H>0$ and have the equations for dark energy and dark matter, and in 3-D where we include the equation for the Hubble function. We also defined dimensionless variables. In Sec. III we explored the parameter space in order to obtain a phase space topology in 2-D with a non-singular high-energy repellor that trajectories can expand from, and a late-time attractor such that trajectories asymptotically tend towards a cosmological constant at late-times. We also defined compact variables to be able to study the dynamics at infinity. We found a high energy repellor and low-energy attractor when $\epsilon=-1$, $10<\mathcal{R}<$ 1, $w_{x}>-1+\mathcal{R}$ and $q=\rho_{*}/\rho_{i}<3/\mathcal{R}$. We then set $\epsilon=-1$ and $\mathcal{R}=0.01$, such that the system only depended on two dimensionless parameters, $w_{x}$ and $q$. In reality, we would expect $10^{-120}<\mathcal{R}<10^{-60}$, however the dynamics is unaffected by setting $\mathcal{R}$ to a larger value, and this helps to have readable phase space plots. As $\mathcal{R}\rightarrow 0$, the attractor $dS_{2+}$ representing the low energy cosmological constant just moves closer to the origin.
@@
In general, two cases for the topology exist: one when $q<3$, and the other when $q>3$. When $q<3$, two separatrices in the 2-D phase space exist, which we present in Sec. IV. Some trajectories expand from the high energy repellor and tend toward the late-time low energy attractor $dS_{2+}$, however some trajectories expand from the high energy repellor and evolve toward a critical point representing a singularity at late-times. We have not explicitly analyzed this, however we expect it to be a big-rip singularity as $\rho\rightarrow\infty$ and $H\rightarrow\infty$, which is caused by the phantom behavior. There are qualitatively interesting expanding models when $q<3$, as there are trajectories that have a decelerated expansion phase followed by a late-time acceleration, asymptotic to the late-time cosmological constant represented by $dS_{2+}$. We then presented the corresponding 3-D dynamics for these qualitatively interesting trajectories. In this case, expanding open and flat models emerge from the the high energy repellor, a fixed point in phase space representing a de Sitter model, and evolve toward the low energy cosmological constant represented by $dS_{2+}$. These all have a decelerated phase followed by a late-time acceleration. Positively curved models are either cyclic, which repeatedly contract and expand through bounces and turnarounds, or they bounce once, contracting from a low energy de Sitter state and expanding toward $dS_{2+}$. The positively curved models of qualitative interest are the subset which go through a quasi-de Sitter bounce at high energy, as they have a decelerated expansion phase followed by a late-time acceleration as they evolve towards $dS_{2+}$.
@@
In Sec. V, we show the phase spaces for $q>3$. In this case, only one separatrix between the high energy repellor and low energy attractor exists in the 2-D phase spaces, and therefore all trajectories avoid a singularity. For certain parameter values, there is a case where the separatrix and zero acceleration curve intersect twice. In this case, all trajectories in the 2-D phase space have a decelerated matter-dominated phase during expansion, followed by a late-time acceleration dominated by dark energy, eventually evolving toward the cosmological constant represented by $dS_{2+}$. Therefore in this subcase, all trajectories represent models which qualitatively correspond to the observed Universe, for which we then presented the 3-D dynamics. The expanding open and flat models are qualitatively interesting, emerging from the high energy de Sitter fixed point and expanding to the late time cosmological constant represented by $dS_{2+}$, with either one or two decelerated phases and a late-time accelerated expansion. Some positively curved models are cyclic, and some are asymptotic in the past to a contracting low-energy de Sitter state, have a bounce, then expand, asymptotically evolving toward the low-energy de Sitter fixed point $dS_{2+}$. The bouncing trajectories which are qualitatively interesting have one or two decelerated phases during expansion followed by a late-time acceleration, some with a quasi-de Sitter transition phase at high energy.
@@
Overall, our analysis shows that for our system of interacting dark energy and dark matter, a parameter range exists such that all trajectories represents singularity-free cosmological models, and can qualitatively match our observed Universe. In particular, trajectories expand from a high energy non-singular fixed point, have a decelerated expansion phase where large scale structure could form, and have a late-time acceleration where trajectories asymptotically tend toward a cosmological constant. In future work, we will quantitatively determine whether the decelerated periods are long enough for large scale structure to form. It will also be interesting to include radiation in the set-up, and analyze if there is any effect on the overall dynamical behavior, however we do not expect it to affect the dynamics if it is produced through a reheating phase, as we found in Paper I Burkmar and Bruni (2023).
Acknowledgments
This work has been supported by UK STFC grants ST/W507738/1 and ST/S000550/1.
@@ Appendix A $q<3$
Here, we show examples of the phase spaces for $q<3$, where $dS_{1+}$ is a repelling node and an improper node. We fix $\epsilon=-1$ and $\mathcal{R}=0.01$ as before, and change $w_{x}$ and $q$. The color scheme of the fixed points and critical points in the 2-D $Z$-$X$ phase spaces are given in Table 4, and the color scheme of the curves are given in Table 5. For the 3-D phase spaces, the color scheme of the fixed points and critical points are given in Table 8, and the color scheme of the curves and surfaces are given in Table 9.
@@ A.1 Repelling Node
Fig. 18 shows an example of the $Z$-$X$ phase space where $q<3$ and $dS_{1+}$ is a repelling node. In this case, the repellor-saddle separatrix and the zero acceleration curve intersect twice. Trajectories that evolve to the left of the two separatrices expand from $dS_{1+}$ toward $S_{2+}$. These trajectories either always accelerate if they never touch the zero acceleration curve, or have a deceleration period if they intersect this curve twice. There is also a limiting case where a trajectory touches the red curve and has a point of zero acceleration, but otherwise accelerates. Trajectories that evolve to the right of the separatrices expand from $dS_{1+}$, and asymptotically tend towards a late-time cosmological constant as they approach $dS_{2+}$. The trajectories that evolve to the right of the separatrices are all of interest, as they all cross the red zero acceleration curve twice and therefore have a decelerated period, followed by a late-time accelerated expansion.
@@
Figure 18: The $Z$-$X$ phase space referred to in Appendix A.1. Here, the parameters are set to $w_{x}=-0.9$ and $q=0.7$. In this case, $dS_{1+}$ is a repelling node and the zero acceleration curve intersects the repellor-saddle separatrix twice. Trajectories to the left of the separatrices expand to $S_{2+}$ which represents a singularity, some with and some without a decelerated period, and those to the right of the separatrices all have a decelerated period, then accelerate as they expand towards $dS_{2+}$. Therefore, the trajectories that evolve to the right of the separatrices are qualitatively of interest.
@@
Fig.s 19 and 20 show trajectories on a first integral surface in the $X$-$Y$-$Z$ phase space generated by fixing initial conditions $X_{0}$ and $Z_{0}$, corresponding to a single trajectory in Fig. 18 that evolves to the right of the separatrices and intersects the zero acceleration curve twice. Fig. 19 shows the trajectories with positive spatial curvature, and Fig. 20 shows expanding flat and open trajectories. As in the main part of the paper, the 3-D phase spaces in the appendices all include a red zero acceleration surface, which corresponds to the zero acceleration curve in the 2-D phase spaces. The trajectories with positive spatial curvature in Fig. 19 all have a bounce. Two Einstein fixed points exist in the phase space; one has saddle stability which the CFS passes through, and the other is a center. Trajectories inside the CFS either bounce once between $dS_{2-}$ and $dS_{2+}$, or are cyclic around the center Einstein fixed point. These trajectories are not of interest, as the bouncing trajectories always accelerate, and the cyclic trajectories have a decelerated phase during expansion, but no late-time acceleration. The trajectories that evolve outside the CFS bounce once and contract from $dS_{2-}$, go through a quasi-de Sitter bounce, then expand towards $dS_{2+}$. These trajectories intersect the zero acceleration surface twice during expansion, meaning they accelerate though the bounce, becoming flatter as they start to expand, then have a decelerated phase where large scale structure could form, and then have a final late-time accelerated expansion as they approach a cosmological constant. These trajectories that evolve outside the CFS are therefore the models of interest.
@@
Figure 19: The $X$-$Y$-$Z$ phase space corresponding to Fig. 18, where $w_{x}=-0.9$ and $q=0.7$. The trajectories are plotted on a first integral surface where $X_{0}=0.07$ and $Z_{0}=0.5$, which corresponds to a trajectory in Fig. 18 that evolves to the right of the separatrices. In this figure, only the trajectories with positive spatial curvature are plotted. Two Einstein fixed points (cyan) exist at $Y=0$ on the zero acceleration surface. Trajectories inside the CFS either bounce once, contracting from $dS_{2-}$ and expanding towards $dS_{2+}$, or are cyclic around an Einstein fixed point. The bouncing models inside the CFS always accelerate, and the cyclic models cross the zero acceleration surface once during expansion, meaning they have a decelerated phase but no late-time acceleration. Trajectories outside the CFS bounce once between $dS_{2-}$ and $dS_{2+}$, and during expansion intersect the zero acceleration surface twice, meaning they have a decelerated phase and a late-time acceleration.
@@
Figure 20: The expanding flat (green) and open (blue) trajectories in the $X$-$Y$-$Z$ phase space corresponding to Fig. 18, where $w_{x}=-0.9$ and $q=0.7$. The trajectories themselves are plotted on a first integral surface where $X_{0}=0.07$ and $Z_{0}=0.5$, which corresponds to a trajectory in Fig. 18 that evolves to the right of the separatrices, expanding to $dS_{2+}$. Both the flat and open trajectories intersect the zero acceleration surface twice, meaning they initially accelerate as they expand from $dS_{1+}$ and $dS_{4+}$, respectively, then have a period of deceleration, and then have a final accelerated phase as they expand toward $dS_{2+}$ at low energy.

### v2

@@ V.2.3 Four Einstein fixed points
Figures 1517 show a first integral surface in the three-dimensional phase space, corresponding to a trajectory in Fig. 9 which intersects the $a^{\prime\prime}=0$ curve four times. Figure 15 shows trajectories with positive spatial curvature, which all have a bounce. One CFS is plotted, which passes through one of the Einstein fixed points, which is a saddle, and loops round the Einstein fixed point above it, which is a center. A subset of trajectories within this CFS are cyclic, and have an initial accelerated expansion followed by a decelerated period, but no late-time acceleration. Bouncing trajectories within this CFS have one decelerated period, and those outside the CFS have two decelerated periods, both with a late-time acceleration. These models all contract from $dS_{2-}$, then accelerate through a bounce and become flatter as they start to expand, with those outside the CFS bouncing during a quasi-de Sitter phase; they then have one or two decelerated expanding phases, followed by a late-time acceleration as they evolve to $dS_{2+}$, therefore these bouncing models can all qualitatively match the observed Universe. The Einstein fixed point close to $X=1$ is a center, and the Einstein fixed point close to $X=0$ is a saddle. We expect this saddle fixed point to form part of another CFS, which would loop around the center Einstein fixed point close to $X=1$; however, we do not plot this separatrix as the numerics breaks down at high energy. We do not expect trajectories within this CFS to be qualitatively interesting; we expect bouncing trajectories to always accelerate, and cyclic trajectories to have no late-time acceleration. Figure 16 shows the view of the three-dimensional phase space through the $Z$-$X$ plane, where it is clearer which trajectory this surface corresponds to in the two-dimensional phase space in Fig. 9.
@@
Figure 15: $X$-$Y$-$Z$ phase space corresponding to Fig. 9, where $w_{x}=-0.95$ and $q=3.1$. The trajectories themselves are plotted on a first integral surface where $X_{0}=0.14$ and $Z_{0}=0.3$, which corresponds to a trajectory in Fig. 9 that intersects the red zero acceleration curve four times. In this figure, only the trajectories with positive spatial curvature are plotted. Four Einstein fixed points (cyan) exist at $Y=0$ on the zero acceleration surface. The only trajectories that are not viable are the cyclic trajectories within the CFS, as during expansion there is no late-time acceleration. All other trajectories bounce once and contract from $dS_{2-}$, then expand towards $dS_{2+}$. Those within the CFS have one decelerated period followed by late-time acceleration, and those outside have two decelerated phases with a late-time acceleration.
@@
Figure 16: Side view of Fig. 15. The cyclic trajectories cross the red zero acceleration surface once during expansion. All other trajectories intersect this surface two or four times during expansion, accelerating when they are above the surface and decelerating when they are below it.
@@
Figure 17 shows the expanding flat and open trajectories. These emerge from $dS_{1+}$ and $dS_{4+}$, respectively, and intersect the $a^{\prime\prime}=0$ surface four times. This means flat and open models have two periods of deceleration and three accelerated phases, including a late-time acceleration as they expand towards $dS_{2+}$.
@@
Figure 17: Expanding flat (green) and open (blue) trajectories in the $X$-$Y$-$Z$ phase space corresponding to Fig. 9, where $w_{x}=-0.95$ and $q=3.1$. The trajectories themselves are plotted on a first integral surface where $X_{0}=0.14$ and $Z_{0}=0.3$, which corresponds to a trajectory in Fig. 9 which intersects the red zero acceleration curve four times. The flat and open trajectories emerge from $dS_{1+}$ and $dS_{4+}$, respectively, and cross the red zero acceleration surface four times, eventually expanding toward $dS_{2+}$ at low energy.
@@ VI Conclusions
In this paper we have studied the dynamics of FLRW models with standard pressureless dark matter interacting with dark energy with a quadratic EoS (2). This EoS introduces two energy scales: $\rho_{\Lambda}$ turns out to be an effective cosmological constant, which plays the role of the asymptotic state of all models evolving toward low energies, and $\rho_{*}>\rho_{\Lambda}$ characterizes the quadratic term. The interaction is also quadratic and characterized by an energy scale $\rho_{i}$, and defined so that energy flows from the dark energy to the dark matter. Given this form of interaction, the dark matter energy density $\rho_{m}$ is always positive, while the dark energy density $\rho_{x}$ remains positive if it was always positive in the past. The quadratic terms are only relevant at very high energies, such that singularities can be avoided.
@@
This work is an extension of Paper I Burkmar and Bruni (2023), where we studied the dynamics of FLRW models containing dark matter, radiation and dark energy with a quadratic EoS, without any interaction term. In Paper I Burkmar and Bruni (2023), we found nonsingular models that had an expanding decelerated phase and a late-time acceleration; however, the decelerated period was lost when the effective cosmological constants were set to realistic energy scales, i.e. when the effective low energy cosmological constant $\rho_{\Lambda}$ is close to the observed value of $\Lambda$, and when the characteristic energy scale of the dark energy $\rho_{*}$ is between Planck and inflationary scales. Here, our aim was twofold. First, to see if including an interaction between the dark components allows for nonsingular models, and second, to see if these nonsingular models could evolve over realistic energy scales, with a decelerated period and late-time accelerated expansion.
@@
In Sec. II we presented the system of equations both in two dimensions, where we take $H>0$ and have the equations for dark energy and dark matter, and in three dimensions where we include the equation for the Hubble function. We also defined dimensionless variables. In Sec. III we explored the parameter space in order to obtain a phase space topology in two dimensions with a nonsingular high energy repellor that trajectories can expand from, and a late-time attractor such that trajectories asymptotically tend towards a cosmological constant at late times. We also defined compact variables to be able to study the dynamics at infinity. We found a high energy repellor and low energy attractor when $\epsilon=-1$, $0<\mathcal{R}<1$, $w_{x}>-1+\mathcal{R}$ and $q=\rho_{*}/\rho_{i}<3/\mathcal{R}$. We then set $\epsilon=-1$ and $\mathcal{R}=0.01$, such that the system only depended on two dimensionless parameters, $w_{x}$ and $q$. In reality, we would expect $10^{-120}<\mathcal{R}<10^{-111}$ if $\rho_{\Lambda}$ is close to the current observed dark energy density Aghanim et al. (2020b); Prat et al. (2022) and $\rho_{*}$ is between inflationary and Planck scales Bass (2015). However, the dynamics are unaffected by setting $\mathcal{R}$ to a larger value, and this helps to have readable phase space plots. As $\mathcal{R}\rightarrow 0$, the attractor $dS_{2+}$ representing the low energy cosmological constant just moves closer to the origin.
@@
In general, two cases for the topology exist: one when $q<3$, and the other when $q>3$. When $q<3$, two separatrices in the two-dimensional phase space exist, which we present in Sec. IV. Some trajectories expand from the high energy repellor and tend toward the late-time low energy attractor $dS_{2+}$; however, some trajectories expand from the high energy repellor and evolve toward a critical point representing a singularity at late times. We have not explicitly analyzed this; however, we expect it to be a big-rip singularity as $\rho\rightarrow\infty$ and $H\rightarrow\infty$, which is caused by the phantom behavior. There are qualitatively interesting expanding models when $q<3$, as there are trajectories that have a decelerated expansion phase followed by a late-time acceleration, asymptotic to the late-time cosmological constant represented by $dS_{2+}$. We then presented the corresponding three-dimensional dynamics for these qualitatively interesting trajectories. In this case, expanding open and flat models emerge from the high energy repellor, a fixed point in phase space representing a de Sitter model, and evolve toward the low energy cosmological constant represented by $dS_{2+}$. These all have a decelerated phase followed by a late-time acceleration.
Positively curved models are either cyclic, which repeatedly contract and expand through bounces and turnarounds, or they bounce once, contracting from a low energy de Sitter state and expanding toward $dS_{2+}$. The positively curved models of qualitative interest are the subset which go through a quasi-de Sitter bounce at high energy, as they have a decelerated expansion phase followed by a late-time acceleration as they evolve towards $dS_{2+}$.
@@
In Sec. V, we show the phase spaces for $q>3$. In this case, only one separatrix between the high energy repellor and low energy attractor exists in the two-dimensional phase spaces, and therefore all trajectories avoid a singularity. For certain parameter values, there is a case where the separatrix and zero acceleration curve intersect twice. In this case, all trajectories in the two-dimensional phase space have a decelerated matter dominated phase during expansion, followed by a late-time acceleration dominated by dark energy, eventually evolving toward the cosmological constant represented by $dS_{2+}$. Therefore in this subcase, all trajectories represent models which qualitatively correspond to the observed Universe, for which we then presented the three-dimensional dynamics. The expanding open and flat models are qualitatively interesting, emerging from the high energy de Sitter fixed point and expanding to the late-time cosmological constant represented by $dS_{2+}$, with either one or two decelerated phases and a late-time accelerated expansion. Some positively curved models are cyclic, and some are asymptotic in the past to a contracting low energy de Sitter state, have a bounce, then expand, asymptotically evolving toward the low energy de Sitter fixed point $dS_{2+}$. The bouncing trajectories which are qualitatively interesting have one or two decelerated phases during expansion followed by a late-time acceleration, some with a quasi-de Sitter transition phase at high energy.
@@
Overall, our analysis shows that for our system of interacting dark energy and dark matter, a parameter range exists such that all trajectories represent singularity-free cosmological models, and are qualitatively consistent with our observed universe. In particular, trajectories expand from a high energy nonsingular fixed point, have a decelerated expansion phase where large scale structure could form, and have a late-time acceleration where trajectories asymptotically tend toward a cosmological constant.
[…]
ACKNOWLEDGEMENTS
This work has been supported by UK STFC Grants No. ST/W507738/1 and No. ST/S000550/1.
DATA AVAILABILITY
The data that support the findings of this article are openly available Aghanim et al. (2020a); Burkmar (2025).
@@ Appendix A $q<3$
Here, we show examples of the phase spaces for $q<3$, where $dS_{1+}$ is a repelling node and an improper node. We fix $\epsilon=-1$ and $\mathcal{R}=0.01$ as before, and change $w_{x}$ and $q$. The color scheme of the fixed points and critical points in the two-dimensional $Z$-$X$ phase spaces are given in Table 4, and the color scheme of the curves are given in Table 5. For the three-dimensional phase spaces, the color scheme of the fixed points and critical points are given in Table 8, and the color scheme of the curves and surfaces are given in Table 9.
@@ A.1 Repelling node
Figure 18 shows an example of the $Z$-$X$ phase space where $q<3$ and $dS_{1+}$ is a repelling node. In this case, the repellor-saddle separatrix and the zero acceleration curve intersect twice. Trajectories that evolve to the left of the two separatrices expand from $dS_{1+}$ toward $S_{2+}$. These trajectories either always accelerate if they never touch the zero acceleration curve, or have a deceleration period if they intersect this curve twice. There is also a limiting case where a trajectory touches the red curve and has a point of zero acceleration, but otherwise accelerates. Trajectories that evolve to the right of the separatrices expand from $dS_{1+}$, and asymptotically tend towards a late-time cosmological constant as they approach $dS_{2+}$. The trajectories that evolve to the right of the separatrices are all of interest, as they all cross the red zero acceleration curve twice and therefore have a decelerated period, followed by a late-time accelerated expansion.
@@
Figure 18: $Z$-$X$ phase space referred to in Appendix A.1. Here, the parameters are set to $w_{x}=-0.9$ and $q=0.7$. In this case, $dS_{1+}$ is a repelling node and the zero acceleration curve intersects the repellor-saddle separatrix twice. Trajectories to the left of the separatrices expand to $S_{2+}$ which represents a singularity, some with and some without a decelerated period, and those to the right of the separatrices all have a decelerated period, then accelerate as they expand towards $dS_{2+}$. Therefore, the trajectories that evolve to the right of the separatrices are qualitatively of interest.
@@
Figures 19 and 20 show trajectories on a first integral surface in the $X$-$Y$-$Z$ phase space generated by fixing initial conditions $X_{0}$ and $Z_{0}$, corresponding to a single trajectory in Fig. 18 that evolves to the right of the separatrices and intersects the zero acceleration curve twice. Figure 19 shows the trajectories with positive spatial curvature, and Fig. 20 shows expanding flat and open trajectories. As in the main part of the paper, the three-dimensional phase spaces in the appendixes all include a red zero acceleration surface, which corresponds to the zero acceleration curve in the two-dimensional phase spaces. The trajectories with positive spatial curvature in Fig. 19 all have a bounce. Two Einstein fixed points exist in the phase space; one has saddle stability which the CFS passes through, and the other is a center. Trajectories inside the CFS either bounce once between $dS_{2-}$ and $dS_{2+}$, or are cyclic around the center Einstein fixed point. These trajectories are not of interest, as the bouncing trajectories always accelerate, and the cyclic trajectories have a decelerated phase during expansion, but no late-time acceleration. The trajectories that evolve outside the CFS bounce once and contract from $dS_{2-}$, go through a quasi-de Sitter bounce, then expand towards $dS_{2+}$. These trajectories intersect the zero acceleration surface twice during expansion, meaning they accelerate though the bounce, becoming flatter as they start to expand, then have a decelerated phase where large scale structure could form, and then have a final late-time accelerated expansion as they approach a cosmological constant. These trajectories that evolve outside the CFS are therefore the models of interest.
@@
Figure 19: $X$-$Y$-$Z$ phase space corresponding to Fig. 18, where $w_{x}=-0.9$ and $q=0.7$. The trajectories are plotted on a first integral surface where $X_{0}=0.07$ and $Z_{0}=0.5$, which corresponds to a trajectory in Fig. 18 that evolves to the right of the separatrices. In this figure, only the trajectories with positive spatial curvature are plotted. Two Einstein fixed points (cyan) exist at $Y=0$ on the zero acceleration surface. Trajectories inside the CFS either bounce once, contracting from $dS_{2-}$ and expanding towards $dS_{2+}$, or are cyclic around an Einstein fixed point. The bouncing models inside the CFS always accelerate, and the cyclic models cross the zero acceleration surface once during expansion, meaning they have a decelerated phase but no late-time acceleration. Trajectories outside the CFS bounce once between $dS_{2-}$ and $dS_{2+}$, and during expansion intersect the zero acceleration surface twice, meaning they have a decelerated phase and a late-time acceleration.
@@
Figure 20: Expanding flat (green) and open (blue) trajectories in the $X$-$Y$-$Z$ phase space corresponding to Fig. 18, where $w_{x}=-0.9$ and $q=0.7$. The trajectories themselves are plotted on a first integral surface where $X_{0}=0.07$ and $Z_{0}=0.5$, which corresponds to a trajectory in Fig. 18 that evolves to the right of the separatrices, expanding to $dS_{2+}$. Both the flat and open trajectories intersect the zero acceleration surface twice, meaning they initially accelerate as they expand from $dS_{1+}$ and $dS_{4+}$, respectively, then have a period of deceleration, and then have a final accelerated phase as they expand toward $dS_{2+}$ at low energy.